# Vision Transformers for Multi-Variable Climate Downscaling: Emulating Regional Climate Models with a Shared Encoder and Multi-Decoder Architecture

### v1

@@ 5 Conclusion
In this study, we have explored multi-variable deep learning approaches for climate downscaling, specifically targeting the GCM-to-RCM emulation task. While most deep learning approaches focus on modeling single variables, we have argued that jointly processing multiple variables can improve both efficiency and predictive performance. Redundant modeling across variables can be avoided, and information from one variable can enhance the prediction of another by capturing inter-variable dependencies. To test this hypothesis, we designed three model configurations using a shared Vision Transformer (ViT)-based encoder-decoder architecture. These configurations explore different strategies for handling multiple variables, ranging from fully independent models to joint encoding with separate decoders. This progression allowed us to assess the trade-offs between variable separation and shared representation, culminating in our proposed Multi Var. ViT (1EMD) model, which combines a shared encoder with dedicated decoders for each variable. To rigorously evaluate our methods, we established a consistent and fair comparison framework. All models share the same architecture, differing only in the components necessary to support multi-variable processing. Training hyperparameters, learning rate scheduling, pre-processing pipelines, and hardware resources were kept identical across all experiments. Our results, based on a 10-year test set of held-out data, demonstrate that the proposed Multi Var. ViT (1EMD) model consistently outperforms all other tested configurations. This suggests that combining shared representations with variable-specific decoders is an effective strategy for multi-variable downscaling. Overall, we have introduced an effective and straightforward method for climate downscaling in a multi-variable context. By modeling variables jointly, our results suggest a substantial gain in both performance and computational efficiency. Several open questions remain, including how many variables can be effectively modeled together, which variables should be selected to maximize performance, and whether more sophisticated loss functions could further enhance the model’s capabilities.

### v2

@@ 5 Conclusion
In this study, we have explored multi-variable deep learning approaches for climate downscaling, specifically targeting the GCM-to-RCM emulation task. While most deep learning approaches focus on modeling single variables, we have argued that jointly processing multiple variables can improve both efficiency and predictive performance. Redundant modeling across variables can be avoided, and information from one variable can enhance the prediction of another by capturing inter-variable dependencies. To test this hypothesis, we evaluated multiple deep learning architectures within a unified experimental framework, including both Vision Transformer (ViT)–based models and a multi-variable U-Net baseline. Across single-variable and multi-variable formulations, these configurations span approaches from fully independent models to joint encoding with shared or variable-specific decoders. This progression enables a systematic assessment of the trade-offs between shared representations and variable specialization, culminating in the proposed Multi Var. ViT (1EMD) model, which combines a shared encoder with dedicated decoders for each variable. To rigorously evaluate our methods, we established a consistent and fair comparison framework. All ViT models share the same architecture, differing only in the components necessary to support multi-variable processing. Training hyperparameters, learning rate scheduling, pre-processing pipelines, and hardware resources were kept identical across all experiments. Our results, evaluated on a 10-year held-out test set, show that multi-variable modeling provides both accuracy and efficiency benefits over single-variable approaches. In particular, the proposed Multi Var. ViT (1EMD) consistently achieves the best performance across all variables, reducing the MSE by 5.5 % on average compared to single-variable ViT. In addition, multi-variable models substantially reduce computational cost, resulting in 29–32% lower inference time per variable compared to single-variable models Overall, we have introduced an effective and straightforward method for climate downscaling in a multi-variable context. By modeling variables jointly, our results reveal a substantial gain in both performance and computational efficiency. Several open questions remain, including how many variables can be effectively modeled together, which variables should be selected to maximize performance, and whether more sophisticated loss functions could further enhance the model’s capabilities.